OpenStudy (anonymous):

Which is 729 written as a power of a number? A. 3 tot he power of 7 B. 3 tot he power of 9 C. 9 to the power of 3 D. 81 tot he power of 9 @imqwerty

OpenStudy (anonymous):

I think it's D

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=9%5E3

imqwerty (imqwerty):

729=3*3*3*3*3*3*3 so 729=3^7 A

OpenStudy (anonymous):

Thx @imqwerty and @misty1212

imqwerty (imqwerty):

np

OpenStudy (anonymous):

@misty1212 is correct. It is C. If you check on a calculator, 9^3 is 729.

imqwerty (imqwerty):

sry sry that is 3^6 so its C

imqwerty (imqwerty):

aww man such a silly mistake